OpenStudy (anonymous):

how do I factor 64 t^4 - 25

OpenStudy (1645323):

set equal to 0 so 64 t^4 -25=0

OpenStudy (anonymous):

I'm still confused, sorry. @1645323

OpenStudy (1645323):

so you add 25 , 64t^4=25

OpenStudy (anonymous):

isn't this just solving for t? I need to factor it. @1645323

OpenStudy (1645323):

you dont need to factor just divide 25 by 64 then so t^4=25/64 and then 4th root 25/64

OpenStudy (accessdenied):

This is a difference of squares. a2−b2=(a−b)(a+b) You should rewrite it like so to see what a and b are. Then factor accordingly! \( \begin{align} 64 t^4 - 25 &= (\color{blue}{8t^2})^2 - (\color{green}{5})^2 & \color{blue}a^2 - \color{green}b^2 \\ &= (\color{blue}{8t^2} - \color{green}5)(\color{blue}{8t^2} + \color{green}5) & (\color{blue}a - \color{green}b)(\color{blue}a + \color{green}b) \end{align} \)

OpenStudy (anonymous):

Yes, I need to factor. Solving for T is incorrect. Thanks anyways.